Prove that Tan^2A/secA+1=1-cosA/cosA                  Trigonometric Identities Based
 
 

Grade:10

1 Answers

Shivam Jain
6 years ago
1 + tan^2A = sec^2ASo tan^2A= sec^2A - 1=(secA + 1)(secA -1) ....(1)Solving LHS...Tan^2A/ secA +1Substituting the value of tan^2A from equation 1..(SecA+1)(secA-1) /(secA+1)= secA - 1= (1/cosA) -1 =1-cosA / cosA=RHSHENCE PROVED